Fermi National Laboratory
 

Accelerator Update for June 6, 2007 - June 8, 2007

 

Wednesday June 6

The day shift began with the Tevatron (TeV) recovering from a quench, with Operations monitoring the Recycler stash (88.8E10), with the Antiproton Source (Pbar) stacking (32.5E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) waiting for the colliding beams.

The TeV access began at 9:03 AM to repair the sector A4 beam valves.

At 10:55 AM, Operations reported that the Recycler lost its stash. Operators investigated, but found no cause for the loss. The investigation continued.

The TeV came out of access at 11:33 AM.

Operations put the TeV through a dry squeeze at 1:11 PM.

Operations put the TeV through a wet squeeze at 2:56 PM.

At 3:02 PM, Operations reported that the Tevatron Electron Lens (TEL2) didn’t have an output current. An expert investigated.

CDF made a 20-minute access into their collision hall beginning at 3:34 PM.

Operations reported at 4:14 PM that the Recycler had stashed antiprotons once after the stash was lost during the day shift.

Operations put the TeV through a dry squeeze at 9:22 PM.

Operations put the TeV into shot setup at 9:43 PM.

Operations established store 5479 at 11:33 PM with an initial luminosity of 137.99E30.

Operations reported at 11:56 PM that the Recycler had stashed three times during the evening shift.

Thursday June 7

The midnight shift began with Operations monitoring store 5479 (122.02E30) and the stash (39.0E10), with Pbar stacking (11.6E10), and with NuMI and MiniBooNE taking beam.

At 3:28 AM, the Booster West Anode Power Supply (WAPS) tripped off. It reset okay, but a Booster RF station (BRF12) tripped off after the WAPS was reset. It reset okay, but Operations reported that BRF18’s output had weakened. Operators reduced the amount of beam to MiniBooNE and contacted an RF expert.

At 6:53 AM, BRF16 tripped off. It reset.

At 7:25 AM, a RF expert arrived to investigate BRF18. Operations turned off BRF18 and lowered Booster’s beam intensity. Booster beam was turned off to disconnect BRF18.

At 7:31 AM, Operations reported a “blip” in the Recycler emittances that lasted about 20 minutes. No beam was lost, but no cause for the emittance problem was found.

Operations reported at 7:54 AM that the Recycler had stashed three times during the midnight shift.

Booster resumed sending beam, at a lower intensity due to BRF18 being off, at 8:08 AM.

At 10:23 AM, Operations turned off Booster and WAPS so RF experts could bring BRF18 back on line. This was completed and Booster resumed sending its normal intensity of beam at 10:54 AM.

The Neutron Therapy Facility (NTF) reported at 11:42 AM that they had treated their last patient for the day.

At 1:58 PM, TEL2 tripped off. It reset.

Operations terminated store 5479 at 2:27 PM.

Operations put the TeV into shot setup at 3:37 PM.

Operations reported at 3:59 PM that the Recycler had stashed two times during the day shift.

At 4:20 PM, Operations turned off NuMI for the three minutes it took to load protons into the TeV.

Operations established store 5480 at 5:17 PM with an initial luminosity of 164.94E30.

At 8:40 PM, Operations reported a vacuum burst in MI-60. Everything reset.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Friday June 8

The midnight shift began with Operations monitoring store 5480 (74.78E30) and the stash (117.6E10), with Pbar stacking (45.0E10), and with NuMI and MiniBooNE taking beam.

At 2:30 AM, a Linac Klystron RF station (KRF3) went out of phase. This happened several times over the next couple of hours. Operators investigated and discovered an errant LCW flow meter. A whack with a hammer and the readjustment of the cavity’s water temperature eventually resolved the problem.

Operations reported at 7:49 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/08/2007    email Fermilab